solve this following quadratic equations by factorisation:
a)x²+7x+12=0
24
b)6x²-5x+1=0
Please help im stuck on these, Thanks

Found 2 solutions by london maths tutor, computerboy2010:
Answer by london maths tutor(243)

a)x²+7x+12=0
(x+3)(x+4) = 0
x = -3 or x = -4
b)6x²-5x+1=0
(6x+1)(x-1) = 0
x = -1/6 or x = 1

Answer by computerboy2010(2)